### v3

# On some aspects of local thermodynamical formalism

Abstract
In 2007, Ye & Zhang introduced a version of local topological entropy. Since their entropy function is, as we show under mild conditions, constant for topologically transitive dynamical systems, we propose to adjust the notion in a way that does not neglect the initial transient part of an orbit. We investigate the properties of this “transient” version, which we call translocal entropy, and compute it in terms of Lyapunov exponents for various dynamical systems. We also investigate how this adjustment affects measure-theoretic local (Brin-Katok) entropy and local pressure functions, generalizing some partial variational principles of Ma & Wen.
Mathematics Subject Classification (2010): Primary: 37B40, Secondary: 37A35.
Keywords. topological entropy, local entropy, Brin-Katok entropy, variational principle, topological pressure, local measure-theoretic pressure.
1 Introduction and motivation
Topological and measure-theoretic entropy can be interpreted as indications of complexity of a map $f:X\to X$, but they provide a global view, without specifying on which subsets of $X$ the map is “more complicated” or “less complicated”. This motivated the study of local versions of entropy and of pressure, as we will discuss presently.
Topological entropy is related to measure-theoretic entropy via the variational principle: If $f$ is a continuous map of a compact metric space $(X,d)$ (an assumption we will make throughout the paper), then $h_{\text{\tiny top}}(f)=\sup_{\mu}h_{\mu}(f)$, where the supremum is taken over all $f$-invariant probability measures $\mu$ on $X$. Measures $\mu$ such that $h_{\text{\tiny top}}(f)=h_{\mu}(f)$ are called measures of maximal entropy or maximal measures. In thermodynamic formalism, one tries to maximize the sum of entropy and energy (represented by an integral over a potential function $\phi:X\to{\mathbb{R}}$), called free energy or measure-theoretic pressure. The topological pressure can then be defined by the variational principle as $P_{top}(f)=\sup_{\mu}\{h_{\mu}+\int_{X}\phi\,d\mu\}$, see e.g. [39, 45].
Local versions of topological entropy were used by Bowen [10] to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$ and by Misiurewicz [34], who proved that vanishing of his local entropy implies existence of a maximal measure. Different versions were used by Newhouse [35], and by Buzzi & Ruette [15, 16], to address questions about the (unique) existence of a measure of maximal entropy. All these notions rely on Bowen balls (also called dynamical balls):
$B_{n}(z;\varepsilon)=\{z^{\prime}\in X:d(f^{j}(z),f^{j}(z^{\prime})<\varepsilon\text{ for all }0\leq j<n\}.$
A local view on measure-theoretic entropy goes back to at least Brin & Katok [12]. In particular, Brin-Katok local entropy $h_{\mu}(f,x)$ (see (7) in Section 3.1 for the definition), when integrated over the space yields the global measure-theoretic entropy:
$\int h_{\mu}(f,x)\,d\mu(x)=h_{\mu}(f).$
Variational principles have been derived for this local version as well, see e.g. [21].
1.1 The Ye & Zhang entropy function and translocal entropy.
In 2007, Ye & Zhang [49] introduced a new version of local entropy, which they called entropy function $h_{\text{\tiny top}}:X\to[0,\infty]$.
Their definition, see (3) below, is the starting point of our paper, but one soon discovers that under mild conditions the entropy function is constant: for any $z\in X$ one has $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for very natural dynamical systems, see Corollary 2.1. It is not straightforward to find maps with non-constant entropy functions, see Section 2.1. The reason is that in the Ye & Zhang approach, only the limit behaviour of the orbit of $z$ counts, not the “transient” behaviour at finite time scales. The entropy function does not see if entropy is created “immediately” in small neighbourhoods of $z$, or only at very distant time scales.
For this reason, we propose an adjustment to Ye & Zhang’s definition, which we call the translocal entropy function and denote it as $h_{\omega}$, that takes into account the time scale needed for Bowen balls to reach unit size. As a result, $h_{\omega}(z)$ depends non-trivially on $z\in X$, and Lyapunov exponents at $z$ play a central role. This interaction is most transparent in smooth one-dimensional maps, see Theorem 2.1, where each point has at most one Lyapunov exponent (because there is only one direction in which derivatives can be taken). As representatives of higher-dimensional maps, we study toral automorphisms $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$ based on a $d\times d$ matrix $A$. Theorem 2.2 shows that its translocal entropy function is determined by the Lyapunov exponents, i.e., the logarithms of the eigenvalues of $A$.
In the second half of the paper, we turn to the local approach to pressure and study the translocal version of pressure. Ma & Wen in [32] obtained a partial variational principle for topological entropy. They noticed (see Theorem 3.2) the relations between topological entropy of a map and its upper and lower measure-theoretic entropies. Using a Carathéodory-like construction, elaborated by Pesin [37], we define a dimensional type of topological pressure $P_{Z}(f,\phi)$ for a continuous map $f:X\to X$ with potential $\phi$, restricted to subsets $Z\subset X$. Fixing a Borel probability measure $\mu$ on $X$ we generalize the Ma & Wen result by means of upper and lower local measure-theoretic pressures, see Theorem 3.3. In Theorem 3.4 we show the same result for translocal version of local measure-theoretic pressures.
1.2 Structure of the paper
Our paper is organized as follows. In Section 2, we introduce the entropy function by Ye & Zhang, and study when it is constant and when not. In Section 2.2, we give our adaptation of the entropy function, and give its basic and more advanced properties in the remaining subsections of Section 2.
Section 3 is devoted to local measure-theoretic approach to pressure. Section 3.1 deals with measure-theoretic versions of local entropy, starting with the Brin-Katok entropy. We discuss (local) topological pressure in Section 3.2 and give a variational principle in Section 3.3. The variational principle of our translocal adaptation is covered in Section 3.4. We finish the paper in Section 4 with a discussion and comparison of various notions of local entropy in the literature.
2 Local topological entropy in sense of Ye & Zhang
Let $f:X\to X$ be a continuous map on a compact metric space $(X,d)$. A set $A$ is called $(n,\varepsilon)$-separated if $d(f^{j}(x),f^{j}(y))>\varepsilon$ for some $0\leq j<n$ and all distinct $x,y\in A$. For a closed subset $K\subset X$, we define the topological entropy [^1] This is in line with the Bowen-Dinaburg definition [9, 18], used in e.g. [38, 41, 42], who were eventually interested in the multifractal properties of a measure-theoretic version of local entropy, in the line of the Brin-Katok entropy which we discuss in Section 3.1. $h_{\text{\tiny top}}(f,K)$ restricted to $K$ as the limit of exponential growth rates of $(n,\varepsilon)$-separated subsets of $K$ as $\varepsilon\to 0$, that is
$h_{\text{\tiny top}}(f,K)=\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K),$ (1)
where $S(n,\varepsilon,K)$ is the maximal cardinality of $(n,\varepsilon)$-separated subsets of $K$.
Can one modify (1) into a pointwise definition of topological entropy by shrinking $K$ to a point? One way of doing this is
$\lim_{\delta\to 0}\inf_{K\in{\mathcal{K}}_{\delta}(x)}h_{\text{\tiny top}}(f,K),$ (2)
where ${\mathcal{K}}_{\delta}(x)$ is the collection of closed neighborhoods of $x$ of diameter $\leq\delta$. A point $x\in X$ is an entropy point if there are arbitrarily small closed neighbourhoods $K$ such that $h_{\text{\tiny top}}(f,K)>0$ and a full entropy point if in addition $h_{\text{\tiny top}}(f,K)=h_{\text{\tiny top}}(f)$, see [49, Section 2]. Hence, only maps with positive entropy can have full entropy points.
A map $f:X\to X$ is topologically exact or locally eventually onto if for every open $U\subset X$, there is $N\in{\mathbb{N}}$ such that $f^{N}(U)=X$. However, if $f$ is locally eventually onto, then there is $N_{K}$ such that $f^{N_{K}}(K)=X$. Notice that $(n-N_{k},\varepsilon)$-separated points $y_{1},y_{2}\in f^{N_{K}}(K)$ determine points $x_{1},x_{2}\in K$, with $y_{1}=f^{N_{K}}(x_{1})$ and $y_{2}=f^{N_{K}}(x_{2})$, such that $x_{1},x_{2}$ are $(n,\varepsilon)$-separated. Therefore $S(n,\varepsilon,K)\geq S(n-N_{k},\varepsilon,f^{N_{K}}(K))$. Thus the quantity in (2) is constant and equal to $h_{\text{\tiny top}}(f)$.
Instead, in [49, Definition 4.1] the local entropy function is defined as
$h_{\text{\tiny top}}(x)=\lim_{\varepsilon\to 0}h_{\text{\tiny top}}(x,\varepsilon),\qquad\text{for~~~}h_{\text{\tiny top}}(x,\varepsilon)=\lim_{\delta\to 0}\inf_{K_{x}\in{\mathcal{K}}_{\delta}(x)}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}).$ (3)
This differs from (2) in the order in which the limits $\varepsilon\to 0$ and $\delta\to 0$ are taken. In general, $h_{\text{\tiny top}}(x)=h_{\text{\tiny top}}(f(x))$, for any point $x\in X$, so the local entropy is constant along forward orbits, see [49, Proposition 4.4(1)]. In fact, we have the following.
Lemma 2.1
The entropy function takes its minimum at transitive points, i.e., points with a dense orbit.
Proof. Fix $\delta>0.$ Choose a point $x\in X$ with a dense orbit and a point $y\in X$ with $K_{y}\in{\mathcal{K}}_{\delta}(y).$ Then, there exists $n_{0}\geq 0$ such that $f^{n_{0}}(x)\in K_{y}$ and a set $K_{x}\in{\mathcal{K}}_{\delta}(x)$ such that $f^{n_{0}}(K_{x})\subset K_{y}$. For any $(n,\varepsilon)$-separated points $y_{1},y_{2}\in f^{n_{0}}(K_{x})$ there exist $(n+n_{0},\varepsilon)$-separated points $x_{1},x_{2}\in K_{x}$ (i.e. $y_{1}=f^{n_{0}}(x_{1})$ and $y_{2}=f^{n_{0}}(x_{2})$). Thus,
$S(n+n_{0},\varepsilon,K_{x})\leq S(n,\varepsilon,f^{n_{0}}(K_{x}))\leq S(n,\varepsilon,K_{y})$
and taking limsup with $n\to\infty$ we get
$\limsup_{n\to\infty}\frac{n+n_{0}}{n}\cdot\frac{1}{n+n_{0}}\log S(n+n_{0},\varepsilon,K_{x})\leq\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{y}).$
Applying (3) and taking the infimum over $K\in{\mathcal{K}}_{\delta}$ on both sides, we obtain
$h_{\text{\tiny top}}(x,\varepsilon)=\inf_{K_{x}\in{\mathcal{K}}_{\delta}(x)}\limsup_{n\to\infty}\frac{1}{n}\log S(n+n_{0},\varepsilon,K_{x})\leq\inf_{K_{y}\in{\mathcal{K}}_{\delta}(y)}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{y})=h_{\text{\tiny top}}(y,\varepsilon),$
and hence $h_{\text{\tiny top}}(x)\leq h_{\text{\tiny top}}(y)$. $\Box$
2.1 When is the entropy function non-constant?
One can argue the merits of local entropy function that is equal to $h_{\text{\tiny top}}(f)$ everywhere. In this case, the behavior near a single orbit suffices to estimate the topological entropy, but on the other hand, such an entropy function cannot indicate any inhomogeneity in where in the system, and how rapidly, chaos is created.
The next theorem shows that the Ye & Zhang local entropy function is indeed constant under mild assumptions. A version of this theorem, for much more specific dynamical systems, can be found in [8].
Topological exactness gives a simple sufficient condition for the entropy function to be constant.
Lemma 2.2
If $f:X\to X$ is topologically exact map, then $h_{\text{\tiny top}}(z)\equiv h_{\text{\tiny top}}(f)$ for any $z\in X$.
Proof. Given a closed neighbourhood $K\owns z$, take $N$ such that $f^{N}(K)\supset f^{N}(\mathring{K})=X$, where $\mathring{K}$ indicates the interior of $K.$ Then $\frac{1}{n}\log S(n,\varepsilon,K)\geq\frac{1}{n}\log S(n-N,\varepsilon,X)=\frac{n-N}{n}\frac{1}{n-N}\log S(n-N,\varepsilon,X)$. So after taking limits $n\to\infty$ and $\varepsilon\to 0$, we find $h_{\text{\tiny top}}(z)\geq h_{\text{\tiny top}}(f)$. $\Box$
Ye & Zhang [49, Theorem 3.7] prove, based on a lemma by Katok, that for every ergodic $f$-invariant measure $\mu$:
$\lim_{\varepsilon\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}\geq h_{\mu}(f).$ (4)
This allows us to derive the following corollary:
Corollary 2.1
Let $f:X\to X$ be a continuous topological transitive map defined on a compact metric space $(X,d)$, and assume that it has at least one measure of maximal entropy. If $h_{\text{\tiny top}}(f)=0$ or if each measure of maximal entropy is fully supported, then $h_{\text{\tiny top}}(x)=h_{\text{\tiny top}}(f)$ for every $x\in X$.
Proof. If $h_{\text{\tiny top}}(f)=0,$ then of course the corollary is true. Since all measure of maximal entropy are fully supported, certainly there is a fully supported ergodic $f$-invariant measure $\mu$ of maximal entropy. Therefore $\mu(K)>0$ for every compact neighbourhood of every $z\in X$. Fix $\delta>0$ and $x\in X.$ Since
$\inf$ $\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\},$
we have
$\lim_{\delta\to 0}$ $\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}.$
Taking limit with $\varepsilon\to 0$ we get
$h_{\text{\tiny top}}(x)$ $=$ $\lim_{\varepsilon\to 0}\lim_{\delta\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{x}):K_{x}\in{\mathcal{K}}_{\delta}(x)~and~\mu(K_{x})>0\}$
$\geq$ $\lim_{\varepsilon\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a Borel set with }\mu(K)>0\}.$
Formula (4) and the assumption that $\mu$ is a measure of maximal entropy yield
$h_{\text{\tiny top}}(f)\geq h_{\text{\tiny top}}(x)\geq h_{\mu}(f)=h_{\text{\tiny top}}(f).$
$\Box$
The assumption that the measures of maximal entropy are fully supported is not entirely automatic, although it is conjectured that on manifolds the measure of maximal entropy (if existent) is automatically fully supported if the map is topologically transitive, see [50] for results supporting this conjecture. In the context of subshifts $(X,\sigma)$ of positive entropy (so $X$ is a Cantor set, not a manifold), Kwietniak et al. [30] presented a counter-example with all of its measures of maximal entropy supported on a proper subsets of $X$. Loosely inspired by their construction, we have the following topologically transitive (and in fact coded) subshift with a non-constant entropy function.
Example 2.1
We construct a coded subshift of $\{0,1,2\}^{\mathbb{Z}}$ on which the local entropy function is not constant. Let $\{w_{k}\}_{k\in{\mathbb{N}}}$ be an enumeration of all words in $\bigcup_{n\geq 1}\{0,1\}^{n}$ such that $|w_{k}|\leq|w_{k+1}|$ for all $k\in{\mathbb{N}}$. Let ${\mathcal{C}}=\{C_{k}\}_{k\in{\mathbb{N}}}$ be the collection of code words, where
$C_{k}=20^{(10+k)!}w_{k}0^{(10+k)!}2.$
A coded shift $X_{{\mathcal{C}}}$ is the shift of which the language consists of subwords of free concatenations of code words, see [31] and [36]. The structure of the code words in $\mathcal{C}$} implies that $22$ is a synchronizing word (i.e., if $u22$ and $22v$ are both allowed words in $X_{{\mathcal{C}}}$, then so is $u22v$), but more importantly, $h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)\geq\log 2$, because the full shift on $\{0,1\}$ is a subshift of $X_{{\mathcal{C}}}$.
Now consider the cylinder set $2.2]=\{z\in\{0,1,2\}^{{\mathbb{Z}}}:z_{-1}=z_{0}=2\}$ and let $z\in[2.2]$ be such that $\operatorname{orb}_{\sigma}(z)$ is dense in $X_{{\mathcal{C}}}$ and $K\subset[2.2]$ is a closed neighbourhood of $z$. From the theory of coded shifts, see [14, Section 3.3], we can compute the exponential growth rate of the number of centered words $x$ of length $2n+1$ and with $x_{-1}=x_{0}=2$. Namely, this is the unique positive solution $h$ of $\sum_{k}e^{-h|C_{k}|}=1$, and it is clear that $h<\log 2$. It follows that $h_{\text{\tiny top}}(z)\leq h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)$. In this case, work of Pavlov [36, Theorem 1.1] implies that its measures of maximal entropy are supported on $\overline{\operatorname{orb}_{\sigma}(z)}\setminus\operatorname{orb}_{\sigma}(z)$, i.e., $\{0,1\}^{\mathbb{Z}}$. On the other hand, for every closed neighbourhood $K$ of $z$, there is $n$ such that $\sigma^{n}(K)\supset[2.2]$, and therefore $h_{\text{\tiny top}}(z)=\log h<h_{\text{\tiny top}}(X_{{\mathcal{C}}},\sigma)=\log 2$.
Without the assumption of transitivity, maps with non-constant entropy are not hard to find. A simple example of a map where $h_{\text{\tiny top}}(z)$ is not constant, and in fact, $\sup_{z}h_{\text{\tiny top}}(z)=\infty$, is given in Figure 1, see [27] for Hölder continuous maps of this type. Maps of this type (i.e., with infinitely many transitive components, each with its own value for local entropy, and infinite entropy altogether), are $C^{0}$-generic on any manifold of dimension $\geq 2$, see [20]. For such examples, the values $E$ that local entropy assumes forms a countable subset of $0,\infty)$. With some more work (but giving up on genericity) one can find smooth examples where $E$ is dense in $0,\infty)$, as well.
Figure 1: An infinite entropy map where $h_{\text{\tiny top}}(x)=\log(2n+1)$ for $x\in(2^{-n},2^{1-n}]$ and $h_{\text{\tiny top}}(0)=\infty$.
2.2 Translocal entropy
Definition 2.1
For $\omega\geq 0$ and $z\in X$, the upper translocal entropy at $z\in X$ is
$\overline{h}_{\omega}(z)=\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,\overline{B(z,e^{-\omega n})}).$
The lower translocal entropy $\underline{h}_{\omega}(z)$ at $z\in X$ is the same with $\liminf$ instead of $\limsup$. If the two are the same, then $h_{\omega}(z):=\overline{h}_{\omega}(z)=\underline{h}_{\omega}(z)$ is the translocal entropy at $z$.
Clearly $\overline{h}_{\omega}(z)$ is decreasing in $\omega$. Compared to the quantity (2), $h_{\text{\tiny top}}(x)$ is obtained by swapping the limit $\lim_{\varepsilon\to 0}$ and $\inf_{K}$.
As a result we get the following.
Lemma 2.3
Let $f:X\to X$ be a continuous map on a compact metric space. For all $x\in X$ and $\omega\geq 0$, we have
$0$ $=$ $\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\inf\{\frac{1}{n}\log S(n,\varepsilon,K_{x}):\delta>0,~K_{x}\in{\mathcal{K}}_{\delta}(x)\}$
$\leq$ $\underline{h}_{\omega}(x)\ \leq\ \overline{h}_{\omega}(x)$
$\leq$ $\inf\{h_{top}(f,K):\delta>0,~K_{x}\in{\mathcal{K}}_{\delta}(x)\}\leq\ h_{\text{\tiny top}}(x).$
Proof. If $K$ is sufficiently small compared to $n$, then $S(n,\varepsilon,K)=1$, so the first equality in (2.3) is obvious. Since $S(n,\varepsilon,K)\geq 1$, we have $\frac{1}{n}\log S(n,\varepsilon,K)\geq 0$ for all $\varepsilon>0$, $n\geq 1$ and closed neighbourhoods $K$ of $x$. Also, $\frac{1}{n}\log S(n,\varepsilon,K)$ is decreasing in $n$ and in $\varepsilon$, but increasing in $K$. Now we prove the second inequality in (2.3). To this aim we fix $n\in\mathbb{N}$ and $\varepsilon>0,$ then there exists $\delta_{n}<2e^{-\omega n}$ such that for any $K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)$ the inclusion $K_{x}\subset B(x,e^{-\omega n})$ holds. Therefore, for any $n\in{\mathbb{N}}$ we get
$\frac{1}{n}\log S(n,\varepsilon,K_{x})\leq\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})).$
Taking infimum over $\delta_{n}>0$ and $K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)$ we obtain
$\inf\left\{\frac{1}{n}\log S(n,\varepsilon,K_{x}):\delta_{n}>0\text{ and }K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)\right\}\leq\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})).$
Finally, taking liminf $n\to\infty$ and limit $\varepsilon\to 0$ we get the second inequality in the statement of the lemma. The third inequality is obvious.
It remains to show the last inequality in (2.3). Fix $\delta>0$ and choose $K\in{\mathcal{K}}_{\delta}(x),$ then there exists $n_{K,\delta}\in\mathbb{N}$ such that for any for any $n\geq n_{K,\delta}$ the inequality $S(n,\varepsilon,B(x,e^{-\omega n})\leq S(n,\varepsilon,K)$ holds. Therefore, for any $n\geq n_{K,\delta}$
$\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\frac{1}{n}\log S(n,\varepsilon,K).$
Thus, taking infimum over $K\in{\mathcal{K}}_{\delta}(x),$ for large enough $n$ we get
$\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\inf\{\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\},$
so
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\}$
and
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\lim_{\delta\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\}.$
Finally, taking limit $\varepsilon\to 0$ we get the last inequality in (2.3). $\Box$
The next example explores the role of Lyapunov exponents in the translocal entropy function. For one-dimensional maps, the Lyapunov exponent is defined as
$\lambda(x)=\lim_{n\to\infty}\frac{1}{n}\log|Df^{n}(x)|,$
provided the derivatives and the limit exist.
Example 2.2
(i) Let $f:{\mathbb{S}}^{1}\to{\mathbb{S}}^{1}$, $x\mapsto 3x\mod 1$. For any $x\in{\mathbb{S}}^{1}$ and $K=[x-\delta,x+\delta]$, we have $f^{n_{0}}(K)={\mathbb{S}}^{1}$ for $n_{0}\geq-\frac{\log 2\delta}{\log 3}$, and $S(n,\varepsilon,K)\geq 3^{n-n_{0}}/\varepsilon$. Therefore $h_{\text{\tiny top}}(f,K)=\log 3=\lambda(x)$ for every $x\in{\mathbb{S}}^{1}$, and so every point is a full entropy point. Regarding $h_{\omega}(z)$, if $K_{n}(z)=\overline{B(z,e^{-\omega n}})$, then after $k=\omega n/\log 3$ iterates, $f^{k}(K_{n}(z))={\mathbb{S}}^{1}$. Therefore
$\frac{1}{n}\log S(n,\varepsilon,K_{n}(z))$ $\sim$ $\frac{1}{n}\log S(n-k,\varepsilon,[0,1])$
$=$ $(1-\frac{\omega}{\log 3})\frac{1}{n-k}\log S(n-k,\varepsilon,[0,1])\to(1-\frac{\omega}{\log 3})h_{\text{\tiny top}}(f).$
This is independent of $z$ because the Lyapunov exponent is the same for every $z$.
(ii) However, if $g:{\mathbb{S}}^{1}\to{\mathbb{S}}^{1}$ is given by
$g(x)=\begin{cases}2x&\text{ if }\ 0\leq x<\frac{1}{2},\\[2.84526pt] 4x-2&\text{ if }\ \frac{1}{2}\leq x<\frac{3}{4},\\[2.84526pt] 4x-3&\text{ if }\ \frac{3}{4}\leq x<1,\end{cases}$
then $h_{\text{\tiny top}}(g,K)=\log 3$, independent of the Lyapunov exponent which is non-constant for this map. Again every point is a full entropy point. But this time, the Lyapunov exponent $\lambda(z)$ varies with the point, and it takes $k\sim\frac{\omega}{\lambda(z)}$ iterates for $K_{n}(z)$ to reach large scale. If the limit $\lambda(z)$ exists, then $h_{\omega}(z)=(1-\frac{\omega}{\lambda(z)})h_{\text{\tiny top}}(f)$.
(iii) The same method shows that $h_{\omega}(z)=0$ for $z=0$ and the Pomeau-Manneville map
$g_{PM}(x)=\begin{cases}\frac{x}{1-x}&\text{ if }\ 0\leq x\leq\frac{1}{2},\\[2.84526pt] 2x-1&\text{ if }\ \frac{1}{2}<x<1.\end{cases}$
For this map, $g_{PM}(\frac{1}{n})=\frac{1}{n-1}$ for $n\geq 2$, so if $K_{n}(0)=[0,e^{-\omega n}]$, then it takes $e^{\omega n}\gg n$ iterates to reach unit size. The above computation gives $h_{\omega}(0)=0$ for all $\omega>0$.
(iv) On the other hand, $h_{\omega}(z)=\log 2$ for $z=0$ and every $\omega\geq 0$ for the map
$g_{\sqrt{}}(x)=\begin{cases}\sqrt{2x}&\text{ if }\ 0\leq x<\frac{1}{2},\\[2.84526pt] 2x-1&\text{ if }\ \frac{1}{2}\leq x<1.\end{cases}$
This time, $g^{k}_{\sqrt{}}(e^{-\omega n})=2^{1-2^{-k}}e^{-\omega n2^{-k}}$, so it takes no more than $k\approx\log n$ iterates to reach large scale. Hence, $h_{\omega}(0)=h_{\text{\tiny top}}(g_{\sqrt{}})=\log 2$ for all $\omega\geq 0$.
Figure 2: The maps $g$, $g_{PM}$ and $g_{\sqrt{}}$.
2.3 Properties of translocal entropy
As the translocal entropy $h_{\omega}(z)=0$ if $f$ is constant on a neighbourhood of $z$, we make the assumption:
In the remainder of this paper, $f:X\to X$ is a local homeomorphism.
Lemma 2.4
The translocal entropy for iterates of the map $f$ satisfies $\underline{h}_{\omega}(f^{R},z)=R\underline{h}_{\omega}(f,z)$ for integers $R\geq 0$ and also for $\underline{h}_{\omega}(f^{-R},z)=R\underline{h}_{\omega}(f,z)$ for integers $R\geq 0$ if $f$ is invertible. The analogous statements hold for $\overline{h}_{\omega}$.
Proof. The proof goes as for the analogous statement on topological entropy. $\Box$
Definition 2.2
We say that $u$ has a super-exponential approach to $v$ if
$\limsup_{k\to\infty}\ -\frac{1}{k}\log d(f^{k}(u),v)=\infty.$
Clearly a periodic point $u$ has a super-exponential approach to every $v\in\operatorname{orb}_{f}(u)$, but the definition refers to other situations too. The following example shows this, but also that super-exponential approach is not a symmetric or transitive notion.
Example 2.3
Let $X=\{0,1\}^{\mathbb{Z}}$ and $\sigma:X\to X$ be the left shift. Let $w=0^{\infty}.0^{\infty}\in X$ the sequence that is constant 0. Let
$v=1^{\infty}.101^{3!-2!}0^{4!-3!}1^{5!-4!}0^{6!-5!}1^{7!-6!}0^{8!-7!}\dots$
Note that the number of iterates $2+\sum_{j=1}^{n-1}j!-(j-1)!=(n-1!)+1$ before the $n$-th block of 0s is reach is negligible compared to the length n!(n1)!=(n1)(n1)!n!-(n-1)!=(n-1)(n-1)! of that block, so $v$ exponentially approaches $w$, but as $w$ is a fixed point, $w$ does not approach $v$. Finally, let
$u=1^{\infty}.1^{2!}v_{1}\dots v_{3!-2!}1^{4!-3!}v_{1}\dots v_{5!-4!}1^{6!-5!}v_{1}\dots v_{7!-6!}1^{8!-7!}\dots$
Then $u$ has super-exponential approach to $v$, but not to $w$.
Proposition 2.1
Let $f:X\to X$ be a Lipschitz continuous map on a metric space. If $u$ has a super-exponential approach to $v$, then $\underline{h}_{\omega}(u)\leq\overline{h}_{\omega}(v)$.
Proof. If $\omega=0$, then the translocal entropy function coincides with the topological entropy, and there is nothing to prove. So assume that $\omega>0$.
Let $L$ be a global Lipschitz constant of $f$; for simplicity, we assume that $\log 2L>1+2\omega$. Choose $\varepsilon>0$ and $R>\log L$ arbitrary and pick $k\in{\mathbb{N}}$ such that $d(f^{k}(u),v)\leq e^{-Rk}<\varepsilon/4$. Then for $n=\lceil Rk/\omega\rceil$ we have
$f^{k}(\overline{B(u,e^{-\omega(n+k)})})\subset B(f^{k}(u),e^{-\omega(n+k)+k\log L})\subset B(v,2e^{-\omega(n+k)+k\log L})\subset\overline{B(v,e^{-\omega(n-k^{\prime})})}$
for $k^{\prime}:=k(\frac{\log 2L}{\omega}-1)>k$.
We can assume that $\operatorname{diam}f^{j}(\overline{B(u,e^{-\omega(n+k)})})<\varepsilon$ for $j\leq k$, so in the first $k$ iterates, no points in $\overline{B(u,e^{-\omega(n+k)})}$ are $\varepsilon$-separated. Therefore
$\frac{1}{n+k}\log S(n+k,\varepsilon,\overline{B(u,e^{-\omega(n+k)})})$ $\leq$ $\frac{1}{n+k}\log S(n,\varepsilon,\overline{B(f^{k}(u),e^{-\omega(n+k)+k\log L})})$
$\leq$ $\frac{1}{n+k}\log\left(S(n,\varepsilon,\overline{B(v,e^{-\omega(n-k^{\prime})})})\right)$
$\leq$ $\frac{1}{n+k}\log\left(S(n-k^{\prime},\varepsilon,\overline{B(v,e^{-\omega(n-k^{\prime})})})\cdot e^{(k^{\prime}+k)h_{\text{\tiny top}}(f)}\right)$
$\leq$ $\frac{1}{n-k^{\prime}}\log S(n-k^{\prime},\varepsilon,\overline{B(v,e^{-\omega(n-k^{\prime})})})+\frac{h_{\text{\tiny top}}(f)\log 2L}{R},$
where in the last line we used $\frac{k^{\prime}+k}{n+k}<\frac{\log 2L}{R}$. Since $h_{\text{\tiny top}}(f)\leq\log L$ for Lipschitz maps and $\varepsilon$ and $R$ are arbitrary, $\underline{h}_{\omega}(u)\leq\overline{h}_{\omega}(v)$, as claimed. $\Box$
Example 2.4
This example shows that no other inequality holds in Proposition 2.1. Let $v$ and $w$ be the one-sided versions of the sequences $v$ and $w$ in Example 2.3, and interpret them as the itineraries of two points $p_{v}\in[0,1]$ and $p_{w}=0$ for the map $g$ in Example 2.2(ii), with respect to the partition $I_{0}=[0,\frac{1}{2})$, $I_{1}=[\frac{1}{2},\frac{3}{4})$ and $I_{2}=[\frac{3}{4},1]$. Then $v$ approaches $w$ super-exponentially, but $\overline{h}_{\omega}(p_{u})=(1-\frac{\omega}{\log 4})\log 3>(1-\frac{\omega}{\log 2})\log 3=h_{\omega}(p_{w})$ for every $\omega\in(0,\log 2]$.
2.4 Translocal entropy for one-dimensional maps
The next theorem is exemplary for the translocal entropy of one-dimensional maps. It holds by and large also for smooth interval maps with critical points, but dealing with the technicalities of distortion control is not the purpose of this paper, so we only state it for expanding circle maps.
Theorem 2.1
Let $f:{\mathbb{S}}^{1}\to{\mathbb{S}}^{1}$ be a $C^{2}$ expanding circle map and $\omega\geq 0$, then for every $z\in{\mathbb{S}}^{1}$
$\overline{h}_{\omega}(z)=\max\{\left(1-\omega/\overline{\lambda}(z)\right)h_{\text{\tiny top}}(f)\ ,\ 0\},$
where $\overline{\lambda}(z)=\limsup_{n\to\infty}\frac{1}{n}\log|Df^{n}(z)|$ is the upper Lyapunov exponent of $z$, and
$\underline{h}_{\omega}(z)=\left(1-\omega/\underline{\lambda}(z)\right)h_{\text{\tiny top}}(f)\},$
where $\underline{\lambda}(z)=\liminf_{n\to\infty}\frac{1}{n}\log|Df^{n}(z)|$ is the lower Lyapunov exponent of $z$.
Proof. If $\omega=0$, then $B(z,e^{-\omega n})$ is independent of $n$, and the topological transitivity gives $\underline{h}_{0}(z)=\overline{h}_{0}(z)=h_{\text{\tiny top}}(f)$ for all $z\in{\mathbb{S}}^{1}$. So let us fix $\omega>0$.
Since $f$ is a $C^{2}$ expanding circle map, there are $c,c^{+},c_{2}>0$ such that for all $x\in{\mathbb{S}}^{1}$ we have expansion: $c^{+}\geq|f^{\prime}(x)|\geq c>1$ and $f^{\prime\prime}(x)\leq c_{2}$. Therefore Adler’s condition: $A(f,x):=|f^{\prime\prime}(x)|/|f^{\prime}(x)|^{2}\leq A:=c_{2}/c^{2}$ holds for all $x\in{\mathbb{S}}^{1}$ (see [1, 11]). This means that Adler’s conditions holds uniformly over all iterates. Indeed, due to a recursive formula $A(f\circ g,x)=(Af,g(x))+\frac{1}{f^{\prime}\circ g(x)}\cdot A(g,x)$, we get
$\frac{|(f^{n})^{\prime\prime}|}{|(f^{n})^{\prime}|^{2}}\leq\frac{Ac}{c-1}\qquad\text{ for all }n\geq 0.$
Fix $\eta:=\frac{c-1}{2cA}$ and abbreviate $K_{n}(z)=\overline{B(z,e^{-\omega n})}$. Let
$k_{n}(z)=\max\{k\geq 0:\operatorname{diam}(f^{k}(K_{n}(z)))\leq\eta\}.$
Take $\xi\in K_{n}(z)$ such that $(f^{k_{n}(z)})^{\prime}(\xi)|K_{n}(z)|=|f^{k_{n}(x)}(K_{n}(z))|\in[\eta/c^{+},\eta]$. Then
$\left|\frac{1}{(f^{k_{n}(z)})^{\prime}(\xi)}-\frac{1}{(f^{k_{n}(z)})^{\prime}(z)}\right|=\left|\int_{z}^{\xi}\left(\frac{1}{(f^{k_{n}(z)})^{\prime}(x)}\right)^{\prime}\,dx\right|=\int_{z}^{\xi}A(f^{k_{n}(z)},x)\,dx\leq\frac{Ac}{c-1}|\xi-z|.$
This gives
$\left|1-\left|\frac{(f^{k_{n}(z)})^{\prime}(\xi)}{(f^{k_{n}(z)})^{\prime}(z)}\right|\right|\leq\frac{Ac}{c-1}\left|(f^{k_{n}(z)})^{\prime}(\xi)\right|\,|\xi-z|\leq\frac{Ac}{c-1}\eta=\frac{1}{2},$
so that $\frac{1}{2}\leq\left|\frac{(f^{k_{n}(z)})^{\prime}(\xi)}{(f^{k_{n}(z)})^{\prime}(z)}\right|\leq\frac{3}{2}$. Recalling that $|K_{n}(z)|=2e^{-\omega n}$, we get $\frac{\eta}{3c^{+}}e^{\omega n}\leq|(f^{k_{n}(z)})^{\prime}(z)|\leq\eta e^{\omega n}$, so that
${\omega}+\frac{1}{n}\log\frac{\eta}{3c^{+}}\leq\frac{k_{n}(z)}{n}\frac{1}{k_{n}(z)}\log|(f^{k_{n}(z)})^{\prime}(z)|\leq{\omega}+\frac{1}{n}\log\eta.$
In the inferior/superior limit, we get $\liminf_{n\to\infty}\frac{k_{n}(z)}{n}=\frac{\omega}{\overline{\lambda}(z)}$ and $\limsup_{n\to\infty}\frac{k_{n}(z)}{n}=\frac{\omega}{\underline{\lambda}(z)}$, respectively.
Let $N_{\eta}$ be such that $f^{N_{\eta}}(J)\supset{\mathbb{S}}^{1}$ for every interval $J$ of length $\eta$. Then for the statement on the $\limsup$, we get
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{n}(z))$ $\geq$ $\limsup_{n\to\infty}\frac{1}{n}\log S\left(n-k_{n}(z)-N_{\eta},\varepsilon,f^{k_{n}(z)+N_{\eta}}(K_{n}(z))\right)$
$=$ $\limsup_{n\to\infty}\frac{n-k_{n}(z)-N_{\eta}}{n}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}$
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}.$
Taking the limit $\varepsilon\to 0$, we obtain $\overline{h}_{\omega}(z)\geq\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
For the other inequality, note that restrictions $f^{j}|_{\overline{B(z,e^{-\omega n})}}$, $1\leq j\leq k_{n}(z)$, are homeomorphisms, and therefore it can separate at most $1/\varepsilon$ points per iterate. Therefore
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{n}(z))$ $\leq$ $\limsup_{n\to\infty}\frac{1}{n}\log\left(S\left(n-k_{n}(z),\varepsilon,f^{k_{n}(z)}(K_{n}(z)\right)\cdot\frac{k_{n}(z)}{\varepsilon}\right)$
$\leq$ $\limsup_{n\to\infty}\frac{n-k_{n}(z)}{n}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}+\frac{1}{n}\log\frac{k_{n}(z)}{\varepsilon}$
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}.\$
Taking the limit $\varepsilon\to 0$, we obtain $\overline{h}_{\omega}(z)\leq\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$, so $\overline{h}_{\omega}(z)=\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
Similarly, for the statement on the $\liminf$, we get
$\liminf_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{n}(z))$ $\geq$ $\liminf_{n\to\infty}\frac{1}{n}\log S(n-k_{n}(z)-N_{\eta},\varepsilon,f^{k_{n}(z)+N_{\eta}}(K_{n}(z))$
$=$ $\liminf_{n\to\infty}\frac{n-k_{n}(z)-N_{\eta}}{n}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}$
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}.$
Taking the limit $\varepsilon\to 0$, we obtain $\underline{h}_{\omega}(z)\geq\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
The other inequality is obtained as follows:
$\liminf_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K_{n}(z))$ $\leq$ $\liminf_{n\to\infty}\frac{1}{n}\log\left(S\left(n-k_{n}(z),\varepsilon,f^{k_{n}(z)}(K_{n}(z))\right)\cdot\frac{k_{n}(z)}{\varepsilon}\right)$
$\leq$ $\liminf_{n\to\infty}\frac{n-k_{n}(z)}{n}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}+\frac{1}{n}\log\frac{k_{n}(z)}{\varepsilon}$
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}.$
Taking the limit $\varepsilon\to 0$, we obtain $\underline{h}_{\omega}(z)\leq\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$, so $\underline{h}_{\omega}(z)=\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$. This finishes the proof. $\Box$
2.5 Translocal entropy for higher-dimensional maps
Lemma 2.5
If $F=f\times g:X\times Y\to X\times Y$ is a Cartesian product and $z=(x,y)$, then $h_{\omega}(F,z)=h_{\omega}(f,x)+h_{\omega}(g,y)$.
Proof. We can take Cartesian product of $(n,\varepsilon)$-separated sets as $(n,\varepsilon)$-separated sets for the Cartesian product $f\times g$, and there are no more efficient sets. Inserting this in the definition of translocal entropy proves the result. $\Box$
Theorem 2.2
Let $A$ be a unimodular integer $d\times d$-matrix and $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$, $x\mapsto Ax\bmod 1$, the affine toral automorphism determined by $A$, then $h_{\omega}(z)=\sum_{\log|\lambda_{i}|\geq\omega}(\log|\lambda_{i}|-\omega)$, where $\lambda_{i}$, $i=1,\dots,d$, are the eigenvalues of the matrix $A$.
Proof. Let $\mu_{1}>\mu_{2}>\dots>\mu_{k}$ be the set of absolute values of eigenvalues and $\kappa_{j}$ is the number of eigenvalues (counted with multiplicity) of absolute value $\mu_{j}$. Each $\mu_{j}$ represents (possibly several) eigenvalues, and the direct sums of their generalized eigenspaces are denoted as $E_{j}$; it is a hyperplane and we let $1\leq d_{j}\leq\dim(E_{j})$ be the size of the largest Jordan block in the Jordan representation of $F_{A}|_{E_{j}}$. If $K_{n,j}(z)$ is the intersection of $B(z,e^{-\omega n})$ and a local plane in the direction of $E_{j}$, then $F_{A}^{k}(K_{n,j}(z))$ reaches unit size after $k_{n}(z)=Cn^{1-d_{j}}e^{n\log\mu_{j}}$ iterates, for some uniform constant $C>0$. Here we assume that $\log\mu_{j}\geq\omega$, because otherwise $F_{A}^{k}(K_{n,j}(z))$ does not reach unit size for $k\leq n$ iterates. Therefore, $S(n,\varepsilon,K_{n,j}(z))\approx S(n-k_{n}(z),\varepsilon,E_{j}(z))$, where $E_{j}(z)$ is a unit size $\kappa_{j}$-dimensional ball centered at $z$ in the direction of $E_{j}$. Taking the logarithm and the limit $n\to\infty$, we obtain $h_{\omega}(z)=\lim_{n\to\infty}\frac{n-k_{n}(z)}{n}\frac{1}{n-k_{n}(z)}\log S(n-k_{n}(z),\varepsilon,E_{j}(z))=\kappa_{j}\max\{\log\mu_{j}-\omega,0\}$. Locally, $F_{A}$ acts as the Cartesian product of $F_{A}|_{E_{1}(z)}\times\dots\times F_{A}|_{E_{k}(z)}$. Therefore $h_{\omega}(z)=\sum_{\log\mu_{j}\geq\omega}(\log\mu_{j}-\omega)=\sum_{\log|\lambda_{i}|\geq\omega}(\log|\lambda_{i}|-\omega)$, as claimed. $\Box$
In fact, if $F$ is a toral automorphism (or Anosov map) that is not necessarily affine, but has bounded distortion, then the formula becomes $\overline{h}_{\omega}(z)=h_{\text{\tiny top}}(f)-\sum_{\log|\overline{\lambda}_{i}(z)|\geq\omega}(\log|\overline{\lambda}_{i}(z)|-\omega)$, where $\overline{\lambda}_{i}$ stands for the $i$-th upper Lyapunov exponent. The same result works for affine hyperbolic horseshoe maps (or non-affine but with bounded distortion). The global product structure of these maps is important to get this outcome, because the formula can be quite different too, as the following example shows.
Example 2.5
Let $\overline{{\mathbb{D}}}$ be the closed unit disk, and let $f:\overline{{\mathbb{D}}}\to\overline{{\mathbb{D}}}$ be given in polar coordinates by $f(r,\phi)\mapsto(r(2-r),3\phi\pmod{2\pi})$. The derivative of $r\mapsto r(2-r)$ at $r=0$ is $2$. Therefore it takes $K_{n}(z):=\overline{B(z,e^{-\omega n})}$ about $\frac{\omega n+\log\varepsilon}{\log 2}$ iterates to reach size $\varepsilon$, and at this point, the $\phi$-component will let $S(n,\varepsilon,K_{n}(z))$ grow, with exponential rate $\log 3$. It follows that $h_{\omega}(0,0)=(1-\frac{\omega}{\log 2})\log 3$.
It is maybe good at this point to make a comparison with the notion of local entropy of $f:X\to X$ used by Buzzi & Ruette [15, 16] and Newhouse [35]. In their definition, an $(n,\varepsilon)$-Bowen ball takes the place of our exponentially small ball $B(z,e^{-\omega n})$. Thus the local entropy is
$h_{\text{\tiny loc}}(f)=\lim_{\varepsilon\to 0}h_{\text{\tiny loc}}(f,\varepsilon)\quad\text{ for }\quad h_{\text{\tiny loc}}(f,\varepsilon)=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\sup_{x\in X}\log r(n,\delta,B_{n}(x,\varepsilon)).$ (6)
In fact, Buzzi uses $(n,\delta)$-spanning sets (the minimal cardinality of which is denoted by $r$ in (6)) instead of $(n,\delta)$-separating sets, but that is of no consequence. Also, by taking the supremum over $x\in X$, $h_{\text{\tiny loc}}(f)$ becomes a quantity independent of the point in space. Buzzi shows that $h_{\text{\tiny loc}}(f)\leq\frac{\dim M}{r}\log\inf_{n}\root n \of{\|Df^{n}\|_{\infty}}$ if $f:M\to M$ is a $C^{r}$-smooth map on a manifold $M$, and hence $h_{\text{\tiny loc}}(f)=0$ for $C^{\infty}$ maps. But if we adjust Example 2.5 to $f(r,\phi)=(r,3\phi\pmod{2\pi})$, then the point $(0,0)$ gives local entropy $\log 3$.
Compared to $h_{\omega}(z)$, for conformal maps (such as the $C^{2}$ expanding circle maps of Theorem 2.1), we find $h_{\text{\tiny loc}}(f)=h_{\omega}(z)=0$ for $\omega=\lambda(z)$, but for non-conformal maps we don’t expect $h_{\text{\tiny loc}}(f)$ and $h_{\omega}(z)=0$ to coincide.
3 Local measure-theoretic entropy and pressure
3.1 Brin-Katok local measure entropy
For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$ and $f$-invariant Borel probability measure $\mu$, Brin & Katok [12] introduced notions of local lower and local upper measure-theoretic entropy w.r.t. $\mu$ as follows
$\begin{cases}{\underline{h}}_{\mu}(f,x):=\lim_{\varepsilon\to 0}\liminf_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\varepsilon)),\\[5.69054pt] {\overline{h}}_{\mu}(f,x):=\lim_{\varepsilon\to 0}\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\varepsilon)).\end{cases}$ (7)
Theorem 3.1 (Main Theorem and Corollary of [12])
For a continuous map $f:X\to X$ on a compact metric space $(X,d)$ and Borel $f$-invariant probability measure $\mu$, the equality ${\underline{h}}_{\mu}(f,x)={\overline{h}}_{\mu}(f,x):=h_{\mu}(f,x)$ holds for $\mu$-a.e. $x\in X$. Moreover, $h_{\mu}(f,x)$ is $f$-invariant and
$\int_{X}h_{\mu}(f,x)\,d\mu(x)=h_{\mu}(f).$
If $\mu$ is ergodic, then additionally $h_{\mu}(f,x)=h_{\mu}(f)$ for $\mu$-a.e. $x$.
Feng & Huang [21, Theorem 1.2(i)] showed in 2012 the following variational principle for the Brin-Katok local entropy. For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$ and a compact non-empty subset $K\subset X$ they obtain
$h_{\text{\tiny top}}^{FH}(f,K)=\sup\{\underline{h}_{\mu}(f):\mu\text{ \color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}is a Borel probability measure and }\mu(K)=1\},$ (8)
where $\underline{h}_{\mu}(f)$ local lower entropy. However, here $h_{\text{\tiny top}}^{FH}(f,K)$ is an entropy-like quantity constructed in a way similar to the construction of $s$-Hausdorff measure and Hausdorff dimension, called Bowen topological entropy, see [21, Section 2.2]. This is based on a Carathéodory-like construction, but a bit more involved than the Bowen topological entropy used in Theorem 3.2 below, and which we discuss in the next section, generalized to pressure.
3.2 Topological pressure and Carathéodory structures
In 2008, Ma & Wen [32] noticed relationship between topological entropy restricted to a set $Z\subset X$ and local measure entropies on $Z$.
Theorem 3.2 (Theorem 1 in [32])
For a continuous map $f:X\to X$ on a compact metric space $(X,d)$, a Borel probability measure $\mu$, a Borel subset $Z\subset X$ and $s\geq 0$ we have:
  1. 1.
    If ${\underline{h}}_{\mu}(f,x)\leq s$ for all $x\in X$, then $h_{\text{\tiny top}}^{B}(f,Z)\leq s$.
  2. 2.
    ${\overline{h}}_{\mu}(f,x)\geq s$ for all $x\in X$ and $\mu(Z)>0$, then $h_{\text{\tiny top}}^{B}(f,Z)\geq s$.
The Bowen entropy $h_{\text{\tiny top}}^{B}$ used in this result is based on Carathéodory construction which we will discuss in this subsection at length, but for topological pressure, so with the inclusion of a potential function $\phi:X\to{\mathbb{R}}$. By setting $\phi\equiv 0$ in the topological pressure defined below, we get $h_{\text{\tiny top}}^{B}(f)$. Wang & Zhang [46] obtain a similar result for upper local entropies $\overline{h}_{\mu}(f,x)$ and packing topological entropies.
The aim of this subsection is to provide a generalization of Theorem 3.2 to topological pressure. Topological pressure, in the spirit of Carathéodory structures elaborated by Pesin [37], for a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$, and restricted to a subset $Z\subset X$. Fix a continuous map (called potential) $\phi:X\to{\mathbb{R}}$, $N\in{\mathbb{N}}$, $r>0$ and $s\geq 0$, define
$M_{Z}(s,r,\phi,N):=\inf\left\{\sum_{j\in J}\exp[-s\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x_{j}))]:Z\subset\bigcup_{j\in J}B_{n_{j}}(x_{j},r);~n_{j}\geq N\right\},$
where the infimum is over all finite or countable covers of $Z$ by Bowen balls $\{B_{n_{j}}(x_{j},r)\}_{j\in J}$ with $n_{j}\geq N$. Here, the Bowen balls are indexed by the elements of a finite or countable set $J$. Using standard arguments (see [37]) one can easily show that for any $N\in{\mathbb{N}}$, the inequality $M_{Z}(s,r,\phi,N+1)\geq M_{Z}(s,r,\phi,N)$ holds. Therefore, there exists a limit
$M_{Z}(s,r,\phi):=\lim_{N\to\infty}M_{Z}(s,r,\phi,N).$
The graph of the function $s\mapsto M_{Z}(s,r,\phi)$ is very similar to the graph of $s$-Hausdorff measure function, i.e., there exists a unique critical parameter $s_{0}$, where the function $s\mapsto M_{Z}(s,r,\phi)$ drops from $\infty$ to 0. Thus, we can define
$M_{Z}(r,\phi):=\sup\{s\geq 0:M_{Z}(s,r,\phi)=\infty\}=\inf\{s\geq 0:M_{Z}(s,r,\phi)=0\}=s_{0}.$
Fix $r_{1}\leq r_{2}$ and consider a cover $\{B_{n_{j}}(x_{j},r_{1})\}_{j\in J}$ of $Z$, with $n_{j}\geq N$. Then $\{B_{n_{j}}(x_{j},r_{2})\}_{j\in J}$ is a cover of $Z$, so $M_{Z}(s,r_{1},\phi)\geq M_{Z}(s,r_{2},\phi)$. Therefore
$M_{Z}(r_{1},\phi)=\inf\{s\geq 0:M_{Z}(s,r_{1},\phi)=0\}\geq\inf\{s\geq 0:M_{Z}(s,r_{2},\phi)=0\}=M_{Z}(r_{2},\phi).$
This proves that the function $r\to M_{Z}(r,\phi)$ is non-increasing, so there exists a limit
$P_{Z}(f,\phi):=\lim_{r\to 0}M_{Z}(r,\phi).$
The quantity $P_{Z}(f,\phi)$ is called the topological pressure of $f$, restricted to $Z$, with respect to the potential $\phi:X\to{\mathbb{R}}$. Basic properties of the topological pressure are presented in the following lemma.
Lemma 3.1
For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$, $r>0$, $s\geq 0$, $N\in{\mathbb{N}}$ and a continuous potential $\phi:X\to{\mathbb{R}}$ we have:
  1. 1.
    If $Z_{1}\subset Z_{2}\subset X$, then $M_{Z_{1}}(s,r,\phi,N)\leq M_{Z_{2}}(s,r,\phi,N)$.
  2. 2.
    If $Z_{1}\subset Z_{2}\subset X$, then $P_{Z_{1}}(f,\phi)\leq P_{Z_{2}}(f,\phi)$.
  3. 3.
    If $Z=\bigcup_{k\in{\mathbb{N}}}Z_{k}$, then $P_{Z}(f,\phi)=\sup\{P_{Z_{k}}(f,\phi):k\in{\mathbb{N}}\}$.
Next, we give a modification of the classical covering lemma (see Section 2.8.4 in [22]).
Lemma 3.2
[Lemma 1 in [32]] Let $f:X\to X$ be a continuous map on a compact metric space. Let $r>0$ and ${\mathcal{B}}(r)=\{B_{n}(x,r):x\in X,n\in{\mathbb{N}}\}$. For any family ${\mathcal{F}}\subset{\mathcal{B}}(r)$ there exists a subfamily $\mathcal{G}$$\subset{\mathcal{F}}$ consisting of disjoint Bowen balls such that
$\bigcup_{B_{n}(x,r)\in{\mathcal{F}}}B_{n}(x,r)\subset\bigcup_{B_{n}(x,r)\in{\mathcal{G}}}B_{n}(x,3\cdot r).$
3.3 Local measure-theoretic pressures
For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$, a Borel probability measure $\mu$ defined on $X$, a continuous potential $\phi:X\to{\mathbb{R}}$ and $x\in X$, we define the upper local measure-theoretic pressure of $f$ at $x$, with respect to $\mu$, by
${\overline{P}}_{\mu}(f,\phi,x):=\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,\varepsilon)))\right).$ (9)
Similarly, we define the lower local measure-theoretic pressure of $f$ at $x$, with respect to $\mu$, by
${\underline{P}}_{\mu}(f,\phi,x):=\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,\varepsilon)))\right).$
Remark 3.1
Notice that local measure-theoretic pressures, calculated with respect to the potential $\phi\equiv 0$, coincide with local measure-theoretic Brin-Katok entropies. Carvalho & Pérez [17] showed that if $\mu$ is ergodic, then ${\overline{P}}_{\mu}(f,\phi,x)={\underline{P}}_{\mu}(f,\phi,x)$ for $\mu$-a.e. $x$, and that the Brin-Katok Theorem 3.1 extends to local pressure:
$\int_{X}{\overline{P}}_{\mu}(f,\phi,x)\,d\mu=P_{\mu}(f,\phi):=h_{\mu}(f)+\int_{X}\phi\,d\mu.$
We have the following partial variational principle, extending Ma & Wen’s result Theorem 3.2.
Theorem 3.3
Given a continuous map $f:X\to X$ on a compact metric space $(X,d)$, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
1) If ${\overline{P}}_{\mu}(f,\phi,x)\leq s$ for all $x\in Z$, then $P_{Z}(f,\phi)\leq s$.
2) If $\mu(Z)>0$ and ${\underline{P}}_{\mu}(f,\phi,x)\geq s$ for all $x\in Z$, then $P_{Z}(f,\phi)\geq s$.
Proof. (1) Fix a Borel subset $Z\subset X$, a Borel probability measure $\mu$, a continuous potential $\phi:X\to{\mathbb{R}}$ and $\varepsilon>0$. Assume that there exists $s\geq 0$ such that ${\overline{P}}_{\mu}(f,\phi,x)\leq s$ for every $x\in Z$. Define a sequence of sets $(Z_{k})_{k\in{\mathbb{N}}}$ by
$Z_{k}:=\left\{x\in Z:\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\leq s+\varepsilon\text{ for all }0<r<\frac{1}{3k}\right\}.$
Then $Z=\bigcup_{k\in{\mathbb{N}}}Z_{k}$. Now we fix $k\in{\mathbb{N}}$ and $r\in(0,\frac{1}{3k})$. The definition of $Z_{k}$ yields that for any $x\in Z_{k}$, there exists a strictly increasing sequence $(n_{j}(x))_{j\in{\mathbb{N}}}$ with
$\sum_{m=0}^{{n_{j}(x)}-1}\phi(f^{m}(x))-\log(\mu(B_{n_{j}(x)}(x,r)))\leq(s+\varepsilon)\cdot n_{j}(x).$
Notice that $Z_{k}$ is contained in the union of elements of the family
${\mathcal{F}}_{N}:=\{B_{n_{j}(x)}(x,r)):x\in Z_{k},~n_{j}(x)\geq N,j\in J\}.$
By Lemma 3.2, there exists a subfamily ${\mathcal{G}}_{N}$ of ${\mathcal{F}}_{N}$, by pairwise disjoint Bowen balls,
${\mathcal{G}}_{N}:=\{B_{n_{i}(x_{i})}(x_{i},r)):x_{i}\in Z_{k},~n_{i}(x)\geq N,i\in I\},$
where $I\subset J$ such that $Z_{k}\subset\bigcup_{i\in I}B_{n_{i}(x_{i})}(x_{i},3\cdot r)$ and
$\mu(B_{n_{i}(x_{i})}(x_{i},r))\geq\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right).$
Therefore,
$M_{Z_{k}}(s+\varepsilon,3\cdot r,\phi,N)$ $\leq$ $\sum_{i\in I}\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right)$
$\leq$ $\sum_{i\in I}\mu(B_{n_{i}(x_{i})}(x_{i},r)))\leq 1,$
where the last inequality follows by the disjointness of Bowen balls in the family ${\mathcal{G}}_{N}$. Passing to the limit $N\to\infty$, we get
$M_{Z_{k}}(s+\varepsilon,3\cdot r,\phi)\leq 1.$
This means that $M_{Z_{k}}(3\cdot r,\phi)\leq s+\varepsilon$. As $r\to 0$, we obtain $P_{Z_{k}}(f,\phi)\leq s+\varepsilon$. Finally, by Lemma 3.1, we obtain
$P_{Z}(f,\phi)=\sup\{P_{Z_{k}}(f,\phi):k\in{\mathbb{N}}\}\leq s+\varepsilon,$
and consequently
$P_{Z}(f,\phi)=\sup\{P_{Z_{k}}(f,\phi):k\in{\mathbb{N}}\}\leq s,$
since $\varepsilon>0$ is arbitrarily small.
(2) Fix $Z\subset X$ such that $\mu(Z)>0$ and $s\geq 0$. Assume that ${\underline{P}}_{\mu}(f,\phi,x)\geq s$ for all $x\in Z$. Fix $\varepsilon>0$ and for any $k\in{\mathbb{N}}$ define
$V_{k}:=\left\{x\in Z:\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for every }0<r<\frac{1}{k}\right\}.$
Since the sequence of sets $(V_{k})_{k\in{\mathbb{N}}}$ increases to $Z$, by continuity of the Borel probability measure $\mu$ we get
$\lim_{k\to\infty}\mu(V_{k})=\mu(Z)>0.$
There exists $k_{0}\in{\mathbb{N}}$ such that $\mu(V_{k_{0}})>\frac{1}{2}\mu(Z)>0$. Now define a sequence of sets $(V_{k_{0},N})_{N\in{\mathbb{N}}}$ by
$V_{k_{0},N}:=\left\{x\in Z:\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for all }n\geq N,0<r<\frac{1}{k_{0}}\right\}.$
Notice that the sequence $(V_{k_{0},N})_{N\in{\mathbb{N}}}$ increases to $V_{k_{0}}$, so by continuity of the Borel probability measure $\mu$, there exists $N_{0}\in{\mathbb{N}}$ such that $\mu(V_{k_{0},N_{0}})>\frac{1}{2}\mu(V_{k_{0}})>0$. By definition of $V_{k_{0},N_{0}}$, for any $x\in V_{k_{0},N_{0}}$, $n_{j}\geq N_{0}$ and $r\in(0,\frac{1}{k_{0}})$, we have
$\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x))-\log(\mu(B_{n_{j}}(x,r)))>(s-\varepsilon)\cdot n_{j},$
so
$\mu(B_{n_{j}}(x,r))<\exp\left[-(s-\varepsilon)\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x))\right].$
Next, consider a countable cover $\mathcal{C}$} of $V_{k_{0},N_{0}}$ by Bowen balls, defined by
${\mathcal{C}}:=\left\{B_{n_{j}}(y_{j},r):y_{j}\in V_{k_{0},N_{0}},~n_{j}\geq N_{0},~r\in(0,\frac{1}{k_{0}}),~B_{n_{j}}(y_{j},r)\cap V_{k_{0},N_{0}}\neq\emptyset,j\in J\right\}.$
Notice that for any $\delta\in(0,\frac{1}{2}\mu(V_{k_{0},N_{0}}))$, there exists a countable family $J_{\delta}$ such that
$M_{V_{k_{0},N_{0}}}(s-\varepsilon,r,\phi,N)$ $\geq$ $-\delta+\sum_{j\in J_{\delta}}\exp\left(-(s-\varepsilon)\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x_{j}))\right)$
$\geq$ $-\delta+\sum_{j\in J_{\delta}}\mu(B_{n_{j}}(y_{j},r))\geq-\delta+\mu(V_{k_{0},N_{0}})\geq\frac{1}{2}\mu(V_{k_{0},N_{0}}).$
Since $V_{k_{0},N_{0}}\subset Z$, for any potential $\phi:X\to{\mathbb{R}}$ and $N\in{\mathbb{N}}$, Lemma 3.1 gives
$M_{Z}(s-\varepsilon,r,\phi,N)\geq M_{V_{k_{0},N_{0}}}(s-\varepsilon,r,\phi,N)\geq\frac{1}{2}\mu(V_{k_{0},N_{0}})>0.$
Passing to the limit $N\to\infty$, we get $M_{Z}(s-\varepsilon,r,\phi)>0$. This means that $M_{Z}(r,\phi)>s-\varepsilon$. Finally, passing to the limit $r\to 0$ and taking into account that $\varepsilon$ is arbitrarily small, we obtain $P_{Z}(f,\phi)\geq s$. The proof is complete. $\Box$
3.4 The translocal version of measure-theoretic local pressure
In this section we introduce a translocal version of the upper measure-theoretic local pressure, which is a modification of ${\overline{P}}_{\mu}(f,\phi,z)$ from (9), in the following way.
$\overline{P}_{\mu,\omega}(f,\phi,z):=\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right),$
and an analogous replacement for lower measure-theoretic local pressure.
$\underline{P}_{\mu,\omega}(f,\phi,z):=\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right).$
The topological translocal pressure itself we will define in the spirit of [37]. That is, define
$M_{f,Z,\omega}(s,\phi,N):=\inf\left\{\sum_{j\in J}\exp[-s\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x_{j}))]:Z\subset\bigcup_{j\in J}B(x_{j},e^{-\omega n_{j}});~n_{j}\geq N\right\},$
where the infimum is over all finite or countable covers of $Z$ by metric balls $\{B(x_{j},e^{-\omega n_{j}})\}_{j\in J}$ with $n_{j}\geq N$. Let
$\overline{M}_{f,Z,\omega}(s,\phi):=\limsup_{N\to\infty}M_{f,Z,\omega}(s,\phi,N),$
$\underline{M}_{f,Z,\omega}(s,\phi):=\liminf_{N\to\infty}M_{f,Z,\omega}(s,\phi,N).$
Lemma 3.3
The graph of the function $s\mapsto\overline{M}_{f,Z,\omega}(s,\phi)$ has a unique critical parameter $s_{0}$, where the function $s\mapsto\overline{M}_{f,Z,\omega}(s,r,\phi)$ drops from $\infty$ to 0. The graph of the function $s\mapsto\underline{M}_{f,Z,\omega}(s,r,\phi)$ has a unique critical parameter $s_{1}$, where the function $s\mapsto\underline{M}_{f,Z,\omega}(s,\phi)$ drops from $\infty$ to 0.
Proof. Proof of the lemma is similar to the proof of Proposition 1.2 in [37]. $\Box$
Thus, we define
$\overline{P}_{Z,\omega}(f,\phi):=\sup\{s\geq 0:\overline{M}_{f,Z,\omega}(s,\phi)=\infty\}=\inf\{s\geq 0:\overline{M}_{f,Z,\omega}(s,\phi)=0\}=s_{0},$
$\underline{P}_{Z,\omega}(f,\phi):=\sup\{s\geq 0:\underline{M}_{f,Z,\omega}(s,\phi)=\infty\}=\inf\{s\geq 0:\underline{M}_{f,Z,\omega}(s,\phi)=0\}=s_{1}.$
A metric space $(X,d)$ is called boundedly compact if all bounded closed subsets of X are compact. In particular ${\mathbb{R}}^{n}$ and Riemannian manifolds (see [26, p.9]) are boundedly compact.
Lemma 3.4
(Vitali Covering Lemma, Theorem 2.1 in [33]). Let $(X,d)$ be a boundedly compact metric space and a family of closed balls ${\mathcal{B}}=\{\overline{B(x,r)}:x\in X,r>0\}$ in X such that
$\sup\{diam(B(x,r)):B(x,r)\in{\mathcal{B}}\}<\infty.$
Then, there is a finite or countable sequence $B(x_{i},r_{i})\in{\mathcal{B}}$, indexed by $i\in I$, of disjoint balls such that
$\bigcup_{B(x,r)\in{\mathcal{B}}}B(x,r)\subset\bigcup_{i\in I}\overline{B(x_{i},5\cdot r_{i})}.$
Properties analogous to those of Lemma 3.1 hold for this version of topological pressure.
Theorem 3.4
Given a continuous map $f:X\to X$ on a compact metric space $(X,d)$, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
1) If ${\overline{P}}_{\mu,\omega}(f,\phi,x)\leq s$ for all $x\in Z$, then $\overline{P}_{Z,\omega}(f,\phi)\leq s$.
2) If $\mu(Z)>0$ and ${\underline{P}}_{\mu,\omega}(f,\phi,x)\geq s$ for all $x\in Z$, then $\underline{P}_{Z,\omega}(f,\phi)\geq s$.
Proof. (1) Fix a Borel subset $Z\subset X$, a Borel probability measure $\mu$, a continuous potential $\phi:X\to{\mathbb{R}}$ and $\varepsilon>0$. Assume that there exists $s\geq 0$ such that ${\overline{P}}_{\mu,\omega}(f,\phi,x)\leq s$ for every $x\in Z$. Therefore, for $x\in Z$, there exists a strictly increasing sequence $(n_{j}(x))_{j\in{\mathbb{N}}}$ with
$\sum_{m=0}^{{n_{j}(x)}-1}\phi(f^{m}(x))-\log(\mu(B(x,e^{-\omega n_{j}(x)})))\leq(s+\varepsilon)\cdot n_{j}(x).$
Notice that $Z$ is contained in the union of elements of the family
${\mathcal{F}}_{N}:=\{B(x_{j},e^{-\omega n_{j}(x_{j})}):x_{j}\in Z,~n_{j}(x_{j})\geq N,j\in J\}.$
Choose $n_{0}$ minimal such that $e^{\omega n_{0}}\geq 5$. By Lemma 3.4 there exists a subfamily ${\mathcal{G}}_{N}$ of ${\mathcal{F}}_{N}$, indexed by elements of a set $I$, by pairwise disjoint metric balls,
${\mathcal{G}}_{N}:=\{B(x_{i},e^{-\omega n_{i}(x_{i})})):x_{i}\in Z,~n_{i}(x_{i})\geq N,i\in I\},$
where $I\subset J$, such that $Z\subset\bigcup_{i\in I}B(x_{i},e^{-\omega(n_{i}(x_{i})-n_{0})}))$ and
$\mu(B(x_{i},e^{-\omega n_{i}(x_{i})})))\geq\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right).$
Therefore,
$M_{Z,\omega}(s+\varepsilon,\phi,N+n_{0})$ $\leq$ $\sum_{i\in I}\exp\left(-(s+\varepsilon)[n_{i}(x_{i})-n_{0}]+\sum_{m=0}^{{n_{i}(x_{i})}-n_{0}-1}\phi(f^{m}(x_{i}))\right)$
$\leq$ $\exp[(s+\varepsilon)\cdot n_{0}]\sum_{i\in I}\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right)$
$\leq$ $\exp[(s+\varepsilon)\cdot n_{0}]\sum_{i\in I}\mu(B(x_{i},e^{-\omega n_{i}(x_{i}))}))\leq\exp[(s+\varepsilon)\cdot n_{0}],$
where the last inequality follows by the disjointness of metric balls in the family ${\mathcal{G}}_{N}$. Taking the $\limsup$ as $N\to\infty$, we get
$\overline{M}_{Z,\omega}(s+\varepsilon,\phi)\leq\exp[(s+\varepsilon)\cdot n_{0}]<\infty$
and therefore
$\overline{P}_{Z,\omega}(f,\phi)\leq s+\varepsilon.$
Since $\varepsilon>0$ is arbitrarily small, $\overline{P}_{Z,\omega}(f,\phi)\leq s$.
(2) Fix $Z\subset X$ such that $\mu(Z)>0$ and $s\geq 0$. Assume that ${\underline{P}}_{\mu,\omega}(f,\phi,x)\geq s$ for all $x\in Z$. Fix $\varepsilon>0$ and for any $N\in{\mathbb{N}}$, define
$V_{N}:=\left\{x\in Z:\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B(x,e^{-\omega n})))\right)\geq s-\varepsilon,\text{ for all }n\geq N\right\}.$
Notice that the sequence $(V_{N})_{N\in{\mathbb{N}}}$ increases to $Z$, so by continuity of the Borel probability measure $\mu$ there exists $N_{0}\in{\mathbb{N}}$ such that $\mu(V_{N_{0}})>\frac{1}{2}\mu(Z)>0$. By definition of $V_{N_{0}}$, for any $x\in V_{N_{0}}$ and $n_{j}\geq N_{0}$, we have
$\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x))-\log(\mu(B(x_{j},e^{-\omega n_{j}})))>(s-\varepsilon)\cdot n_{j},$
so
$\mu(B(x,e^{-\omega n_{j}})))<\exp\left[-(s-\varepsilon)\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x))\right].$
Next, consider a countable cover $\mathcal{C}$} of $V_{N_{0}}$ by metric balls, defined by
${\mathcal{C}}:=\left\{B(y_{j},e^{-\omega n_{j}}):y_{j}\in V_{N_{0}},~n_{j}\geq N_{0},~B(y_{j},e^{-\omega n_{j}})\cap V_{N_{0}}\neq\emptyset,j\in J\right\}.$
Notice that for any $\delta\in(0,\frac{1}{2}\mu(V_{N_{0}}))$, there exists a countable family $J_{\delta}$ such that
$M_{V_{N_{0}},\omega}(s-\varepsilon,\phi,N)$ $\geq$ $-\delta+\sum_{j\in J_{\delta}}\exp\left(-(s-\varepsilon)\cdot n_{j}+\sum_{m=0}^{n_{j}-1}\phi(f^{m}(x_{j}))\right)$
$\geq$ $-\delta+\sum_{j\in J_{\delta}}\mu(B(x_{j},e^{-\omega n_{j}})))\geq-\delta+\mu(V_{N_{0}})\geq\frac{1}{2}\mu(V_{N_{0}}).$
Since $V_{k_{0},N_{0}}\subset Z$, for any potential $\phi:X\to{\mathbb{R}}$ and $N\in{\mathbb{N}}$, we get
$M_{Z,\omega}(s-\varepsilon,\phi,N)\geq M_{V_{N_{0}}}(s-\varepsilon,\phi,N)\geq\frac{1}{2}\mu(V_{N_{0}})>0.$
Taking the $\liminf$ as $N\to\infty$, we get $\underline{M}_{Z,\omega}(s-\varepsilon,\phi)>0$. This means that $\underline{M}_{Z,\omega}(\phi)>s-\varepsilon$. Finally taking into account that $\varepsilon$ is arbitrarily small, we obtain $\underline{P}_{Z,\omega}(f,\phi)\geq s$. The proof is complete. $\Box$
4 Discussion on local entropy
In the literature on dynamical systems there are several attempts to define local entropy. Bowen [10] introduced the notion of $\varepsilon$-entropy to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$. Another notion of local entropy (called conditional topological entropy) was given by Misiurewicz [34], who proved that vanishing of his local entropy implies existence of maximal measure. Finding (uniqueness of) measures of maximal entropy was also the reason for Newhouse [35] and Buzzi & Ruette [15, 16] to define a version the local entropy $h_{\text{\tiny loc}}(f)$, see (6). Buzzi & Ruette [16] proved that for $C^{1}$ maps $f:[0,1]\to[0,1]$ with critical set $\operatorname{Crit}(f)=\{x\in[0,1]:f|_{U}\text{ is not monotone on each neighborhood }U\owns x\}$, if $h_{\text{\tiny top}}(f)>h_{\text{\tiny top}}(f,\operatorname{Crit}(f))+h_{\text{\tiny loc}}(f)$, then $f$ admits a measure of maximal entropy.
4.1 Orbit complexity
Brudno [13] introduced the notion of orbit complexity for a homeomorphism $f:X\to X$ of a compact space $X$ and related it to measure-theoretic entropy and topological entropy. In Brudno’s approach the orbit complexity is a measure of the amount of information that is necessary to describe the orbit. The complexity $K(x,f)$ of the orbit of $x\in X$ is defined in the spirit of asymptotically optimal functions in the sense of Kolmogorov; it indicates the limit proportion of $n$-prefixes of the symbolic itinerary of $x$ needed to reconstruct the whole $n$-prefix. In his Theorem 3.1, which slightly predates what we state as the Brin-Katok Theorem 3.1 in Section 3.1, Brudno proved the following result: If an $f$-invariant measure $\mu$ is ergodic, then [^2] Entropy is computed with respect to logarithmic base $b$ where $b$ is the cardinality of the alphabet used for the symbolic itineraries. $K(x,f)=h_{\mu}(f)$ for $\mu$-a.e. $x\in X$. Theorem 3.2 in [13] says that for any point $x\in X$ one has $K(x,f)\leq h_{f}(x)$, where $h_{f}(x):=\sup\{h_{\mu}(f):\mu\text{ is a weak accumulation point of }(\frac{1}{n}\sum_{k=0}^{n-1}\delta_{f^{k}(x)})_{n\geq 1}\}$ is the entropy of Kamae introduced in [29].
The results of Brudno were extended by White [47], who proved that for any ergodic $f$-invariant measure $\mu$, $\mu$-a.e. orbit has a limiting complexity equal to $h_{\text{\tiny top}}(f)$ Also, if $f:X\to X$ is minimal, then the set of points with orbit upper complexity equal to the topological entropy is a residual set. If $f:X\to X$ has a unique measure $\mu_{0}$ of maximal entropy, then any point with lower complexity equal to $h_{\text{\tiny top}}(f)$, is generic for the measure $\mu_{0}$. This is likely to hold also if each measure of maximal entropy is fully supported as in Corollary 2.1.
Another extension of results of Brudno [13] one can find in the paper by Galatolo [23], who considered a dynamical system given by a continuous map on a separable metric space, and introduced a notion of computable structure on a metric space with a new notion of orbit complexity with respect to this structure. For compact spaces, Galatolo’s definition coincides with the definition given by Brudno.
A different way of using single orbits to estimate entropy goes back to time series analysis, and is known as Takens’ estimator in applied dynamics and fractal dimension estimation, see e.g. [40, 43]. Viewing orbits $Y:=(f^{n}(x_{0}))_{n\geq 0}$ for $x_{0}\in X$ as time series, one can define $k\times\dim(X)$-dimensional reconstruction vectors $(x_{n},x_{n+1},\dots,x_{n+k-1})$ and estimate the probability ${\mathbb{P}}_{\varepsilon}^{k}(Y)$ of finding two such vectors at mutual distance $\leq\varepsilon$. The entropy is the logarithmic growth rate of this quantity as $\varepsilon\to 0$: $h(Y)=\lim_{\varepsilon\to 0}\lim_{k\to\infty}\frac{1}{k}\log{\mathbb{P}}^{k}_{\varepsilon}(Y)$. Takens’ estimator is a maximum likehood estimator of this quantity, which is also applicable to arbitrary time series, i.e., not just for orbits only.
4.2 Entropy pairs and tuples
In 1993, Blanchard [6] introduced the notion of entropy pairs, in order to localize “where” in the system entropy is generated. This was generalized to entropy $n$-tuples and entropy sets by Huang & Ye [28], Dou, Ye & Zhang [19] and Blanchard & Huang [7].
Definition 4.1
Let $f:X\to X$ be a continuous map on a compact space $X$. A set $E\subset X$ is an entropy set if $E\geq 2$, and if for each open cover ${\mathcal{U}}=\{U_{1},\dots,U_{n}\}$ of $X$ with the property that for each $i$ there is $x\in E\setminus\overline{U_{i}}$, the topological entropy $h_{\text{\tiny top}}(f,{\mathcal{U}})$ (in the sense of Adler-Konheim-McAndrew [2]) of the cover is positive. An entropy set of $n$ points is an entropy $n$-tuple, so an entropy pair is an entropy $2$-tuple.
Dou, Ye & Zhang [19] characterize maximal entropy sets, and prove that $E$ is an entropy set if and only if each $n$-tuple in $E$ with at least two distinct point is an entropy $n$-tuple.
The survey paper of Glasner & Ye [25] is largely on the relation between the topological and ergodic approach to dynamical systems from the global point of view. In particular, they discuss topological and measure entropy pairs and $n$-tuples and present the relation between the topological entropy of open covers and the measure-theoretic entropy of finer partitions. The second part of the survey is on local properties of dynamical systems. They introduce $n$-tuples for an invariant measure, and establish their relation to topological entropy $n$-tuples. Glasner & Ye [25] discuss in detail the notion of entropy set, in order to determine where the entropy is concentrated, including the result of Dou, Ye and Zhang [19, Corollary 3.3], namely that any topological dynamical system with positive entropy admits an entropy set with infinitely many points. Moreover, for any topological dynamical system, there exists a compact countable subset such that its $(n,\varepsilon)$-separated set entropy is equal to the entropy of the whole system, see [49].
Glasner & Ye discuss notions of completely positive entropy and uniform positive entropy, introduced by Blanchard [5] in a topological context: A topological dynamical system has completely positive entropy if every non-trivial factor has positive topological entropy, and it has uniform positive entropy if the topological entropy of every non-dense finite open cover is positive. Uniform positive entropy implies weak mixing as well as completely positive entropy, which in turn implies the existence of a fully supported invariant measure.
4.3 Neutralized entropy
The paper by Garcia-Ramos & Li [24] is also a survey about recent developments in the local entropy theory for topological dynamical systems; they cover similar material as Glasner & Ye, but for general (continuous) group actions.
In 1992, Thieullen [44] introduced a local form of entropy which he called $\alpha$-entropy and which is close to our approach of translocal entropy. In his definition, the $\varepsilon$ in the Bowen $(n,\varepsilon)$-balls is replaced by an $n$-dependent quantity $e^{-\alpha n}$, but unlike our translocal entropy, he still uses Bowen balls. Ben Ovadia & Rodriguez-Hertz [4] introduced neutralized (local) entropy as
${\mathcal{E}}_{\mu}(x):=\lim_{\alpha\to 0}\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,e^{-\alpha n})).$ (10)
This is the $\alpha$-entropy where eventually the limit $\alpha\to 0$ is taken. Ben Ovadia & Rodriguez-Hertz study the role it can play in hyperbolic dynamics and compare it with the Brin-Katok local entropy, see (7).
They show that the neutralized local entropy for $C^{1+\beta}$ diffeomorphism $f:M\to M$ of a compact closed manifold $M$, calculated with respect to an $f$-invariant probability measure $\mu$, coincides with Brin-Katok local entropy almost everywhere.
In a recent paper [48], Yang, Chen & Zhou, extending the variational principle (8) of Feng & Huang [21], obtained a variational principle for neutralized entropy of a continuous map $f:X\to X$ on a compact metric space. For a non-empty compact set $K\subset X$, they prove the equality
[…]
Although neutralized entropy ${\mathcal{E}}_{\mu}(x)$ uses similar components as translocal entropy $h_{\omega}(x)$, they are hard to compare because the former depends on the measure $\mu$ and the latter on the parameter $\omega$.
Acknowledgments The first author was partially supported by the inner Łódź University Grant 11/IDUB/DOS/2021. This research was finalized during the visit of the second author to the Jagiellonian University funded by the program Excellence Initiative – Research University at the Jagiellonian University in Kraków. Both authors are very grateful for the careful reading and the many due corrections by the anonymous referees.